OpenStudy (anonymous):

if sinA/2 and cosB/2 are the roots of px^2 + qx + r -0 and A= pie/3 p + 2q + 4r =0 p+ 2q = 4r

OpenStudy (anonymous):

Put x = 1/2 in the equation: (Because sinA\2 = sin pi/6 = 1/2) \[\frac{p}{4} + \frac{q}{2} + r = 0\] \[p + 2q + 4r = 0\]

OpenStudy (anonymous):

for p+ 2q = 4r i think one of p and q must be 0

OpenStudy (anonymous):